# CalibFormer: A Transformer-based Automatic LiDAR-Camera Calibration Network

### 26 Nov 2023 v1

@@ I Introduction
Some works [9, 10, 11] have attempted to achieve targetless transformation matrices, but they require hand-crafted features and may not work well in different environments. Recently, deep learning techniques have been widely used and demonstrated the superiority of automatic feature engineering. For example, several methods [12, 13, 14] use deep learning models to estimate the 6-DoF transformation between camera and LiDAR coordinates. Most deep learning-based methods utilize RGB images and depth images obtained through point cloud projection – often miscalibrated – as inputs. The processing pipeline primarily involves three steps: feature extraction, feature matching, and calibration parameter regression. However, deep learning-based calibration methods face several important design issues. For example, the balance between calibration accuracy and computation efficiency is hard to achieve. It is also challenging to accurately correlate corresponding features from different modalities due to the disparate physical characteristics and operational principles of sensors, yielding data of varying dimensions, qualities, and types. Furthermore, not all correlated features contribute equally to calibration, making it imperative to extract feature correlations with higher contributions.
@@ II-B Targetless Methods
Regnet [12] was the first to adopt a deep learning approach. It extracts and matches features using a network before regressing the calibration parameters. CalibNet [24] incorporates geometric information by introducing a 3D spatial transformer layer into the model. CalibRCNN [13] combines CNN with LSTM and adds pose constraints between consecutive frames to improve the calibration accuracy. LCCNet [14] utilizes a cost volume to compute the correlation between features from different sensors. In sensor fusion perception tasks such as object detection [1, 25, 26], the transformer’s cross-attention mechanism is used to achieve soft-association between different modalities. Nevertheless, the application of the transformer in calibration methods is rare.
In contrast to these methods, our approach employs a deep aggregation module to obtain high-resolution feature maps from the LiDAR and camera. Furthermore, we incorporate a multi-head correlation module to compute inter-sensor correlations in a more fine-grained manner. We also utilize a transformer architecture to process correlations and estimate calibration parameters effectively.
Figure 1: The overview of our proposed method for camera and LiDAR calibration. Firstly, we project the LiDAR point cloud onto the image plane, generating a miscalibrated LiDAR image using the initial extrinsic parameter $\mathbf{T}_{init}$ and the camera matrix $\mathbf{K}$. Our network takes both RGB images and LiDAR images as inputs. After extracting fine-grained features, we employ a multi-head correlation module and a transformer architecture to obtain a 6-DoF transformation $\mathbf{T}_{pred}$ representing the deviation between the initial extrinsic parameter $\mathbf{T}_{init}$ and the accurate extrinsic parameter $\mathbf{T}_{LC}$.
@@ III-B Fine-grained Feature Extraction
The feature extraction network consists of two symmetric branches for extracting features from the RGB image $X_{rgb}$ and LiDAR image $X_{lidar}$, respectively. In the image branch, we use a pre-trained ResNet-18 [27] model, excluding the final fully connected layer, to extract image features. Meanwhile, the LiDAR branch has a similar structure to the image branch, except that the input channel of the first convolutional layer is replaced with two channels to accommodate the LiDAR image $X_{lidar}$.
We use an enhanced version of the Deep Layer Aggregation (DLA) [15] to fuse multi-scale features with different receptive fields, as illustrated in Fig. 2. Unlike the original DLA [28], it has more skip connections between feature maps of different scales. Additionally, to dynamically adjust the receptive field based on the image, the convolutional layers in the upsampling module of the DLA are replaced with deformable convolution layers [29]. We apply this enhanced DLA on both the image and LiDAR branches separately, resulting in high-resolution feature maps $F_{lidar}$ and $F_{rgb}$, which are used for subsequent calibration.
We additionally employ an rgb-guided query initialization module to extract features $F_{query}$ from RGB images. This module consists of a ResNet-18 and a global average pooling layer, employed for generating the initial pose query.
@@ III-C Multi-head Feature Matching
After separately extracting fine-grained features from the input data, a feature matching module is used to calculate the correlation between the misaligned features of the two modalities caused by the use of the noisy initial parameter $\mathbf{T}_{init}$ during LiDAR point projection in Section III-A. To compute the correlation between features in a fine-grained manner, we apply a multi-head correlation module.
Firstly, the LiDAR feature map $F_{lidar}$ and camera image feature map $F_{rgb}$ are flattened into queries $Q$ and keys $K$, both with a dimension of $d_{k}$. Then, we compute the dot product between the query $Q$ and the key $K$. After that, we divide the result by $\sqrt{d_{k}}$ to obtain the correlation weights. The computation process can be expressed as:
@@
Given that the initial deviation of extrinsic parameters falls within a certain range, the offset of the LiDAR and RGB image features is also expected to be within a certain range. Specifically, for an image with a resolution of $1241\times 376$, if the translation and rotation deviations of the calibration parameters are within $0.5$\mathrm{m}$$ and $5$\mathrm{\SIUnitSymbolDegree}$$ respectively, the projected point cloud deviation is approximately within $100$ pixels. To leverage this characteristic, we apply a window-based approach, in which only the correlation weights within a particular window are utilized for calibration. Given a window size $d$, the correlation weights between the RGB feature $F_{rgb}(p_{i})$ and the LiDAR feature $F_{lidar}(p_{j})$ are only considered if $|p_{i}-p_{j}|_{\infty}\leq d$, where $p_{i}$ and $p_{j}$ are the 2D positions of RGB feature map $F_{rgb}$ and LiDAR feature map $F_{lidar}$, respectively. Therefore, we can obtain the correlation feature map $F_{corr}\in\mathbb{R}^{((2d+1)^{2}\times n)\times H\times W}$, where $H$ and $W$ denote the height and width of the feature map, respectively, and $n$ denotes the number of heads.
@@ III-D Transformer-based Parameter Regression
We employ a transformer architecture to extract calibration parameters from the correlation feature map $F_{corr}$. For the correlation feature $F_{corr}$, we first increase its dimension to $d_{k}$ by densely connected convolutional layers [30]. Subsequently, it is flattened and employed as input $Q$, $K$, and $V$ for the transformer encoder. Given the high resolution of the correlation feature map, we utilize a Swin Transformer [16] encoder to mitigate computational complexity. We employ a Transformer [17] decoder to estimate calibration parameters. The query $Q$ is derived from the initial pose query $F_{query}$, while the key $K$ and value $V$ are obtained from the encoder’s output. The positions are embedded as position encoding using a multi-layer perceptron (MLP) and element-wise added to the key features. Finally, a feed-forward network (FFN) is utilized to regress the pose information, and then we obtain the translation parameters and rotation parameters separately. The translation parameter is represented by a $1\times 3$ vector $\mathbf{t}_{pred}$, and the rotation parameter is represented by a $1\times 4$ quaternion $\mathbf{q}_{pred}$.
@@ IV-B Implementing Details
Training details: The original resolution of the image is padded to $1280\times 384$ and subsequently resized to $512\times 256$ as the input for the model. The upsampling rate for feature maps is set to 4×4$\times$4 ×, the window size in the multi-head correlation module is set to $4$, the number of layers in the Swin Transformer encoder is set to $2$, and the number of layers in the Transformer decoder is set to $6$. We train the network using the Adam optimizer with a learning rate of $5e^{-4}$ for $500$ epochs and a batch size of $256$. Latency is measured on an NVIDIA RTX 3060 GPU.
Evaluation Metrics: The calibration results are evaluated by translation and rotation parameters. For the translation component, we separately record mean absolute errors in the X, Y, and Z directions. Regarding the rotation component, although it can be represented using quaternions, we opt to convert it into Euler angles to provide a more intuitive assessment of rotational error. Similarly, we separately record mean absolute errors of roll, pitch, and yaw.
@@ IV-C Quantitative Results
To compare performance with different methods and follow their experimental setups, we introduce two sets of initial deviations, (±$(\pm 0.5$\mathrm{m}$,\pm 5$\mathrm{\SIUnitSymbolDegree}$)$ and (±$(\pm 0.25$\mathrm{m}$,\pm 10$\mathrm{\SIUnitSymbolDegree}$)$, and evaluate the network’s performance for each. The results, as presented in Table I, indicate that for both initial deviation settings, our proposed method outperforms other approaches.
@@
Ours $\mathbf{0.8751}$ $\mathbf{0.7594}$ $\mathbf{0.6027}$ $\mathbf{1.2633}$ $\mathbf{0.0562}$ $\mathbf{0.0249}$ $\mathbf{0.1041}$ $\mathbf{0.0395}$ $27.79$
@@ IV-D Ablation Studies
Transformer Architecture: We investigate the impact of using a transformer architecture to process correlation features on calibration performance. As a comparison, we substituted a fully connected layer for the transformer architecture. As shown in Table II, our approach yields a moderate increase in latency, resulting in a $29.2\%$ improvement in translation performance and a $25.1\%$ improvement in rotation performance. Furthermore, if only the Transformer decoder is utilized without the Swin Transformer encoder, the calibration results deteriorate. This outcome underscores the contribution of both the transformer encoder and decoder in locating and leveraging more valuable correlation features.
Upsampling: We compare the impact of aggregating the features of different layers, i.e., different upsampling rates. As shown in Table III, we observe that using a higher upsampling rate leads to better results. This is because a higher-resolution output feature map allows for a more detailed representation of the input data, potentially enabling the network to capture more fine-grained features and patterns. However, it is important to consider that using a higher upsampling rate also increases the computational complexity and memory requirements of the network, involving there may be a trade-off between performance and efficiency. As the upsampling rate increases to 8×8$\times$8 ×, the performance improvement becomes marginal compared to that of 4×4$\times$4 ×, while the computational overhead is much higher. Hence, 4×4$\times$4 × is a more appropriate choice.
@@ IV-F Generalization Validation
Ours $\mathbf{1.7314}$ $\mathbf{1.0382}$ $\mathbf{0.7043}$ $\mathbf{1.5685}$ $\mathbf{0.1319}$ $\mathbf{0.0380}$ $0.1556$ $\mathbf{0.0479}$
The KITTI odometry dataset primarily consists of the “2011_09_30” and “2011_10_03” sequences from the KITTI raw dataset, along with a small part of the “2011_09_26” sequence. To evaluate the network’s generalization capability, we train the network on the KITTI odometry dataset and evaluate it on the “2011_09_26” sequence of KITTI raw dataset with unfamiliar scenes. Similarly, the initial deviation is set to (±$(\pm 0.5$\mathrm{m}$,\pm 5$\mathrm{\SIUnitSymbolDegree}$)$. The results shown in Table IV indicate that due to variations in the scene, the performance of our network on the KITTI raw dataset is worse compared to the results on the KITTI odometry dataset. Nevertheless, it still achieves a mean translation error of $1.7314$\mathrm{cm}$$ and a mean rotation error of $0.1319$\mathrm{\SIUnitSymbolDegree}$$. Compared to LCCNet [14], our method demonstrates a $30.4\%$ improvement in translation performance and a $31.9\%$ improvement in rotation performance. The test results on the KITTI raw dataset indicate that our network maintains strong performance across different scenes, showcasing robust generalization capabilities.
@@ V Conclusion
In this paper, we propose an end-to-end calibration network for estimating the 6-DoF rigid body transformation between the LiDAR and the camera, which are important sensor combinations in autonomous driving systems for perception. Our network consists of three main parts: feature extraction module, feature matching module, and transformer regression module. The fine-grained feature extraction module employs DLA to aggregate multi-layer features to get a fine-grained feature representation. In the multi-head feature matching module, we use a multi-head correlation module to calculate the correlation between two modalities in a fine-grained manner. In the transformer-based parameter regression module, we employ both Swin Transformer and Transformer for encoding and decoding, resulting in accurate translation and rotation parameters. In order to solve the problem of insufficient training samples, we introduce random deviations to the extrinsic parameters to augment the training data. Our network can achieve an absolute error of $0.8751$\mathrm{cm}$$ in translation and $0.0562$\mathrm{\SIUnitSymbolDegree}$$ in rotation, with initial miscalibrations up to ±$\pm 0.5$\mathrm{m}$$ in translation and ±$\pm 5$\mathrm{\SIUnitSymbolDegree}$$ in rotation, outperforming other state-of-the-art methods. Despite the increased computational cost, the latency remains within a reasonable range.

### 17 Mar 2024 v2

@@ I Introduction
Some works [9, 10, 11] have attempted to obtain the transformation matrix using targetless methods, but they require hand-crafted features and may not work well in different environments. Recently, deep learning techniques have been widely used and demonstrated the superiority of automatic feature engineering. For example, several methods [12, 13, 14] use deep learning models to estimate the 6-DoF transformation between camera and LiDAR coordinates. Most deep learning-based methods utilize RGB images and depth images obtained through point cloud projection — often miscalibrated — as inputs. The processing pipeline primarily involves three steps: feature extraction, feature matching, and calibration parameter regression. However, deep learning-based calibration methods face several important design issues. For example, the balance between calibration accuracy and computation efficiency is hard to achieve. It is also challenging to accurately correlate corresponding features from different modalities due to the disparate physical characteristics and operational principles of sensors, yielding data of varying dimensions, qualities, and types. Furthermore, not all correlated features contribute equally to calibration, making it imperative to extract feature correlations with higher contributions.
@@ II-B Targetless Methods
Regnet [12] was the first to adopt a deep learning approach. It extracts and matches features using a network before regressing the calibration parameters. CalibNet [24] incorporates geometric information by introducing a 3D spatial transformer layer into the model. CalibRCNN [13] combines CNN with LSTM and adds pose constraints between consecutive frames to improve the calibration accuracy. LCCNet [14] utilizes the cost volume to compute the correlation between features from different sensors. Moreover, in sensor fusion perception tasks such as object detection [1, 25, 26], the transformer’s cross-attention mechanism is used to achieve soft association between different modalities. Nevertheless, the application of the transformer in calibration methods is rare.
In contrast to these methods, our approach employs a deep layer aggregation module to obtain high-resolution feature maps from the LiDAR and camera. Furthermore, we incorporate a multi-head correlation module to compute inter-sensor correlations in a more fine-grained manner. We also utilize a transformer architecture to process correlations and estimate calibration parameters effectively.
Figure 1: The overview of our proposed method for camera and LiDAR calibration. Firstly, we project the LiDAR point cloud onto the image plane, generating a miscalibrated LiDAR image using the initial extrinsic parameter $\mathbf{T}_{init}$ and the camera matrix $\mathbf{K}$. Our network takes both camera images and LiDAR images as inputs. After extracting fine-grained features, we employ a multi-head correlation module and a transformer architecture to obtain a 6-DoF transformation $\mathbf{T}_{pred}$ representing the deviation between the initial extrinsic parameter $\mathbf{T}_{init}$ and the accurate extrinsic parameter $\mathbf{T}_{LC}$.
@@ III-B Fine-grained Feature Extraction
The feature extraction network consists of two symmetric branches for extracting features from the camera image $X_{cam}$ and LiDAR image $X_{lidar}$, respectively. In the camera branch, we use a pre-trained ResNet-18 [27] model to extract image features. Meanwhile, the LiDAR branch has a similar structure to the image branch, except that the number of input channels for the first convolutional layer is adjusted to two to accommodate the LiDAR image $X_{lidar}$.
We use an enhanced version of the Deep Layer Aggregation (DLA) [15] to fuse multi-scale features with different receptive fields, as illustrated in Fig. 2. Unlike the original DLA [28], it has more skip connections between feature maps of different scales. Additionally, to dynamically adjust the receptive field based on the image, the convolutional layers in the upsampling module of the DLA are replaced with deformable convolution layers [29]. We apply this enhanced DLA on both the camera and LiDAR branches separately, resulting in high-resolution feature maps $F_{lidar}$ and $F_{cam}$, which are used for subsequent calibration.
We additionally employ a camera-guided query initialization module to extract features $F_{query}$ from camera images. This module consists of a ResNet-18 and a global average pooling layer, employed for generating the initial pose query.
@@ III-C Multi-head Feature Matching
After separately extracting fine-grained features from the input data, a feature matching module is used to calculate the correlation between the misaligned features of the two modalities caused by the noisy initial parameter $\mathbf{T}_{init}$ during LiDAR point projection in Section III-A. To compute the correlation between features in a fine-grained manner, we apply a multi-head correlation module.
Firstly, the LiDAR feature map $F_{lidar}$ and camera feature map $F_{cam}$ are flattened into queries $Q$ and keys $K$, both with a dimension of $d_{k}$. Then, we compute the dot product between the query $Q$ and the key $K$. After that, we divide the result by $\sqrt{d_{k}}$ to obtain the correlation weights. The computation process can be expressed as:
@@
Given that the initial deviation of extrinsic parameters falls within a certain range, the offset of the LiDAR and camera image features is also expected to be within a certain range. Specifically, for an image with a resolution of $1241\times 376$ from the KITTI dataset, if the translation and rotation deviations of the calibration parameters are within $0.5$\mathrm{m}$$ and $5$\mathrm{\SIUnitSymbolDegree}$$ respectively, the projected point cloud deviation is approximately within $100$ pixels. To leverage this characteristic, we apply a window-based approach, in which only the correlation weights within a particular window are utilized for calibration. Given a window size $d$, the correlation weights between the camera feature $F_{cam}(p_{i})$ and the LiDAR feature $F_{lidar}(p_{j})$ are only considered if $|p_{i}-p_{j}|_{\infty}\leq d$, where $p_{i}$ and $p_{j}$ are the 2D positions of camera feature map $F_{cam}$ and LiDAR feature map $F_{lidar}$, respectively. Therefore, we can obtain the correlation feature map $F_{corr}\in\mathbb{R}^{((2d+1)^{2}\times n)\times H\times W}$, where $H$ and $W$ denote the height and width of the feature map, respectively, and $n$ denotes the number of heads.
@@ III-D Transformer-based Parameter Regression
We employ a transformer architecture to extract calibration parameters from the correlation feature map $F_{corr}$. For the correlation feature $F_{corr}$, we first increase its dimension to $d_{k}$ by densely connected convolutional layers [30]. Subsequently, it is flattened and employed as input $Q$, $K$, and $V$ for the transformer encoder. Given the high resolution of the correlation feature map, we utilize a Swin Transformer encoder [16] to mitigate computational complexity. We employ a Transformer decoder [17] to estimate calibration parameters. The query $Q$ is derived from the initial pose query $F_{query}$, while the key $K$ and value $V$ are obtained from the encoder’s output. The positions are embedded as position encoding using a multi-layer perceptron (MLP) and element-wise added to the key features. Finally, a feed-forward network (FFN) is utilized to regress the pose information, and then we obtain the translation parameters and rotation parameters separately. The translation parameter is represented by a $1\times 3$ vector $\mathbf{t}_{pred}$, and the rotation parameter is represented by a $1\times 4$ quaternion $\mathbf{q}_{pred}$.
@@ IV-B Implementing Details
Training details: The original resolution of the image is padded to $1280\times 384$ and subsequently resized to $512\times 256$ as the input for the model. The upsampling rate for feature maps is set to $4$. The window size in the multi-head correlation module is set to $4$. Furthermore, the Swin Transformer encoder is configured with $2$ layers, while the Transformer decoder comprises $6$ layers. We train the network using the Adam optimizer [32] with a learning rate of $5e^{-4}$ for $500$ epochs and a batch size of $256$. Latency is measured on an NVIDIA RTX 3060 GPU.
Evaluation Metrics: The calibration results are evaluated by translation and rotation parameters. For the translation component, we separately record mean absolute errors in the X, Y, and Z directions. Regarding the rotation component, although it can be represented using quaternions, we opt to convert it into Euler angles to provide a more intuitive assessment of rotation error. Similarly, we separately record mean absolute errors of roll, pitch, and yaw.
@@ IV-C Quantitative Results
To compare performance with different methods, we follow their experimental setups, introducing two sets of initial deviations, (±$(\pm 0.5$\mathrm{m}$,\pm 5$\mathrm{\SIUnitSymbolDegree}$)$ and (±$(\pm 0.25$\mathrm{m}$,\pm 10$\mathrm{\SIUnitSymbolDegree}$)$, and evaluate the network’s performance for each. The results, as presented in Table I, indicate that for both initial deviation settings, our proposed method outperforms other approaches.
@@
w/ all modules $\mathbf{0.8751}$ $\mathbf{0.7594}$ $\mathbf{0.6027}$ $\mathbf{1.2633}$ $\mathbf{0.0562}$ $\mathbf{0.0249}$ $\mathbf{0.1041}$ $\mathbf{0.0395}$ $27.79$
@@ IV-D Ablation Studies
Transformer Architecture: We investigate the impact of using a transformer architecture to process correlation features on calibration performance. As a comparison, we substitute a fully connected layer for the transformer architecture. As shown in Table II, our approach yields a moderate increase in latency, resulting in a $29.2\%$ improvement in translation performance and a $25.1\%$ improvement in rotation performance. Furthermore, if only the Transformer decoder is utilized without the Swin Transformer encoder, the calibration results deteriorate. This outcome underscores the contribution of both the transformer encoder and decoder in locating and leveraging more valuable correlation features.
Upsampling: We compare the impact of aggregating the features of different layers, i.e., different upsampling rates. As shown in Table III, we observe that using a higher upsampling rate leads to better results. This is because a higher-resolution output feature map allows for a more detailed representation of the input data, potentially enabling the network to capture more fine-grained features and patterns. However, it is important to consider that using a higher upsampling rate also increases the computational complexity and memory requirements of the network, involving there may be a trade-off between performance and efficiency. As the upsampling rate increases to $8$, the performance improvement becomes marginal compared to that of $4$, while the computational overhead is much higher. Hence, a 4×4$\times$4 × upsampling rate is a more appropriate choice.
@@ IV-F Generalization Validation
Ours $\mathbf{1.1037}$ $\mathbf{1.0382}$ $\mathbf{0.7043}$ $\mathbf{1.5685}$ $\mathbf{0.0805}$ $\mathbf{0.0380}$ $0.1556$ $\mathbf{0.0479}$
The KITTI odometry dataset primarily consists of the “2011_09_30” and “2011_10_03” sequences from the KITTI raw dataset, along with a small part of the “2011_09_26” sequence. To evaluate our model’s generalization capability, we train the network using the KITTI odometry dataset and subsequently evaluate its performance on the “2011_09_26” sequence of KITTI raw dataset, which includes unfamiliar scenes. Similarly, the initial deviation is set to (±$(\pm 0.5$\mathrm{m}$,\pm 5$\mathrm{\SIUnitSymbolDegree}$)$. The results shown in Table IV indicate that due to variations in the scene, the performance of our network on the KITTI raw dataset is worse compared to the results on the KITTI odometry dataset. Nevertheless, it still achieves a mean translation error of $1.1037$\mathrm{cm}$$ and a mean rotation error of $0.0805$\mathrm{\SIUnitSymbolDegree}$$. In comparison to LCCNet [14], our method demonstrates a $55.7\%$ improvement in translation performance and a $58.4\%$ improvement in rotation performance. The test results on the KITTI raw dataset underscore the robustness of our network, demonstrating consistent and strong performance across diverse scenes.
@@ V Conclusion
In this paper, we propose an end-to-end calibration network for estimating the 6-DoF rigid body transformation between the LiDAR and the camera, which are important sensor combinations in autonomous driving systems for perception. Our network consists of three main parts: feature extraction module, feature matching module, and transformer regression module. The fine-grained feature extraction module employs DLA to aggregate multi-layer features to get a high-resolution feature representation. In the multi-head feature matching module, we use a multi-head correlation module to calculate the correlation between two modalities in a fine-grained manner. In the transformer-based parameter regression module, we employ both Swin Transformer and Transformer for encoding and decoding, resulting in accurate translation and rotation parameters. In order to solve the problem of insufficient training samples, we introduce random deviations to the extrinsic parameters to augment the training data. Our network can achieve an absolute error of $0.8751$\mathrm{cm}$$ in translation and $0.0562$\mathrm{\SIUnitSymbolDegree}$$ in rotation, with initial miscalibrations up to ±$\pm 0.5$\mathrm{m}$$ in translation and ±$\pm 5$\mathrm{\SIUnitSymbolDegree}$$ in rotation, outperforming other state-of-the-art methods. Despite the increase in computational cost, the latency remains within a reasonable range.